SAGE MINDTAP WULL HOW Fly 10t? Search this course rk 5: Confidence intervals anment Homework 5: Confidence Intervals Asment S comfidence interval and margin of error
Answer #1

a. Here Pshort = 0.1913

b. For 95% CI, z value is 1.96 as P(-1.96<z<1.96)=0.95

So Margin of error is E = 2* = 1.96 * 0.1913 * 0.8087 -= 0.0606 162 V n

So CI is CI = PIE=0.1913 + 0.0606 = (0.1307,0.2519)

c. For Margin of Error to be 0.05, we will find n using formula of E

E = 2* P9 n

So n= ()2 *p*9= .06 )2 +0.1913 # 0.8087 = 237.72 = 238